OpenStudy (anonymous):

Cameron has been eating 2 dollar menu burgers every week(7 days). At that rate, how many hamburgers will he in 4 weeks ?

OpenStudy (danjs):

4 times the amount he will eat in 1 week

OpenStudy (danjs):

4 weeks worth

OpenStudy (anonymous):

good work

OpenStudy (anonymous):

I want to know how makes the steps for understand more ! somebody can explain me ?

OpenStudy (dan815):

he is eating 7 burgers a week

OpenStudy (dan815):

each is 2 dollars

OpenStudy (dan815):

and for 4 weeks

OpenStudy (dan815):

2*7*4=56

OpenStudy (danjs):

eats 2 hambugers each week eats 4 hamburgers in 2 weeks eats 6 hamburgers in 3 weeks eats 8 hamburgers in 4 weeks

OpenStudy (dan815):

oh

OpenStudy (danjs):

each week he adds 2 more hamburgers to his total

OpenStudy (dan815):

i misinterpreted the question

OpenStudy (anonymous):

2 burgers per week * 4 weeks = 8 burgers

OpenStudy (anonymous):

i have to make fractions is gonna be like this 2/7 = x/4 ?

OpenStudy (danjs):

\[\frac{ 4~weeks }{ 1}*\frac{ 2~hamburgers }{ 1~week } = \]

OpenStudy (dan815):

okay

OpenStudy (dan815):

2/7=x/(4*7)

OpenStudy (dan815):

2/7 =x/28 2/7=8/28

OpenStudy (anonymous):

So is 4/x= 2/7 thats equal 14 , im no sure !

OpenStudy (dan815):

for 1 week there are 7 days

OpenStudy (dan815):

he is eating 2 every 7 days

OpenStudy (dan815):

2/7, for 4 weeks there are 4*7 days = 28 days how many does he eat in 28 days? x/28

OpenStudy (dan815):

the ratio is the same

OpenStudy (dan815):

2/7=x/28 solve for x

OpenStudy (anonymous):

im sorry i'm so confused i dont know how put the steps

OpenStudy (perl):

the question is a bit ambiguous is he eating two 'dollar menu burgers' a week, or is eating one 'two dollar menu' burgers a week.

OpenStudy (perl):

lets say its the former, he eats two burgers a week.

OpenStudy (perl):